OpenStudy (anonymous):

Can someone solve for x and y using the method: Substitution y = 1/2x - 6 2x + 6y = 19 Thanks.(:

OpenStudy (campbell_st):

take the 1st equation and substitute it into the 2nd equation for y so 3x + 6(1/2 x - 6) = 19 distribute, collect like terms and solve for x

OpenStudy (anonymous):

I don't understand how to solve to x... and did you even get y.?

OpenStudy (mathstudent55):

To solve a system of equations by substitution you need to follow these three steps: 1. Solve one equation for one variable. 2. Substitute what that variable is equal to in the other equation and solve for the other variable. Now one variable is known. 3. Substitute the known variable into one of the two original equations, and solve for the other varrable. Let's do it here, one step at a time: Step 1. Since the first equation is already solved for y, step 1 is done, y = (1/2)x - 6 Step 2. Substitute what y is equal to in the other equation: 2x + 6y = 19, but y = (1/2)x - 6, so 2x + 6( (1/2)x - 6) = 19 Distribute the 6 on the left side: 2x + 3x - 36 = 19 Combine 2x and 3x: 5x - 36 = 19 Add 36 to both sides: 5x = 55 Divide both sides by 5: x = 11 Step 3. Substitute x in one of the two original equations: y = (1/2)x - 6 y = (1/2)11 - 6 y = 11/2 - 6 Use LCD of 2 to subtract fractions: y = 11/2 - 12/2 y = -1/2 Solution: x = 11, y = -1/2

OpenStudy (anonymous):

You are my best friend dude. This helps me out so well. Thank you.

OpenStudy (mathstudent55):

I suggest you go over this answer carefully a few times. This is not just the solution to this system of equations but a guide as to how to solve a system using substitution. In fact, refer to it when you need a refresher. You're very welcome.

OpenStudy (anonymous):

Yeah, I copied and pasted it to my notes in my computer. Loll.(:

OpenStudy (anonymous):

yeah! that the answer ;)